# Indirect detection of the QCD axion

### v1

@@ 1 Introduction
The strong-CP problem can be addressed by introducing a new global symmetry into the Standard Model (SM) of particle physics, as proposed by Peccei and Quinn (PQ) [4, 5]. This solution predicts the existence of the QCD axion, a pseudo-scalar particle [6, 7], which could serve as dark matter (DM) if produced through non-thermal processes that ensure the axion remains non-relativistic by the time of recombination [8, 9, 10]. An active search for the QCD axion and related axion-like particles is ongoing, see Refs. [11, 12, 13] for reviews, though detection is challenging due to the axion’s weak interactions with SM particles. Most experimental efforts focus on the axion-photon coupling, , which modifies Maxwell’s equations in the presence of axions [14, 15, 16, 17, 18, 19, 20]. This coupling motivates various terrestrial experiments [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] and astrophysical searches for axion-photon conversion in regions with strong magnetic fields, such as those surrounding neutron stars (NSs) [40, 41, 42, 43, 44, 1].
@@ Acknowledgments
The author thanks all of his collaborators that participated in the papers presented: Prakamya Agrawal, Madeleine Edenton, Thomas Edwards, Bradley Johnson, Bradley Kavanagh, David J. E. Marsh, Scott Ransom, Jordan E. Shroyer, Liam Walters, Christoph Weniger, and Sam Witte. The author also thanks the Green Bank Observatory for providing the observing time on the GBT (awards GBT22A-067 and GBT23A-245). The author acknowledges the support of the National Natural Science Foundation of China (NSFC) through Grant No. 12350610240 “Astrophysical Axion Laboratories”.

### v2

@@ 1 Introduction
The strong-CP problem can be addressed by introducing a new global symmetry into the Standard Model (SM) of particle physics, as proposed by Peccei and Quinn (PQ) [4, 5]. This solution predicts the existence of the QCD axion, a pseudo-scalar particle [6, 7], which could serve as dark matter (DM) if produced through non-thermal processes that ensure the axion remains non-relativistic by the time of recombination [8, 9, 10]. An active search for the QCD axion and related axion-like particles is ongoing, see Refs. [11, 12, 13, 14] for reviews, though detection is challenging due to the axion’s weak interactions with SM particles. Most experimental efforts focus on the axion-photon coupling, , which modifies Maxwell’s equations in the presence of axions [15, 16, 17, 18, 19, 20, 21]. This coupling motivates various terrestrial experiments [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41], searches in helioscopes [42, 43, 44], and astrophysical searches for axion-photon conversion in regions with strong magnetic fields, such as those surrounding neutron stars (NSs) [45, 46, 47, 48, 49, 1].
@@ Acknowledgments
The authors thank their collaborators that helped with the development of the ideas presented in their papers: Prakamya Agrawal, Madeleine Edenton, Scott Ransom, Christoph Weniger, and Sam Witte. The authors thank the Green Bank Observatory for providing the observing time on the GBT (awards GBT22A-067 and GBT23A-245). LV acknowledges the support of the National Natural Science Foundation of China (NSFC) through Grant No. 12350610240 “Astrophysical Axion Laboratories”. BJK acknowledges support from the DMpheno2lab Project (Project PID2022-139494NB-I00 financed by MCIN /AEI /10.13039/501100011033 / FEDER, UE). D.J.E.M. is supported by an Ernest Rutherford Fellowship from the STFC, Grant No. ST/T004037/1. This article is based on the work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).